OpenStudy (anonymous):

Let f(x)=ln[((x^6)(x+3)^8) ((x^2)+7)^6] find f'(x)

OpenStudy (anonymous):

start by using the properties of the log first \[6\ln(x)+8\ln(x+3)-6\ln(x^2+7)\] then taking the derivative should be easy

OpenStudy (anonymous):

so then would you just take the derivative of the insides of the parts?

OpenStudy (anonymous):

take the derivative piece by piece

OpenStudy (anonymous):

for example the derivative of \(6\ln(x)\) is \(\frac{6}{x}\) etc